A 13.02 gram sample of cobalt is heated in the presence of excess iodine. A metal iodide is formed with a mass of 97.12 g. Determine the empirical formula of the metal iodide.
1
Expert's answer
2020-12-09T06:34:43-0500

Solution:

Molar mass of Co 58.33 g/mol

Molar mass of I 126,90 g/mol

Mass I = 97.12 g – 13.02 g = 84.01 g

n Co = 13.02g/58.33g/mol = 0.22 mol

n I = 84.01g/126.90g/mol = 0.66 mol

Since atoms combine in the same ratio that moles do, we divide all of the numbers of moles by the smallest number to put everything into lowest terms:

Co 0.22 I0.66 → Co 0.22/0.22 I0.66/0.22 → CoI3

Answer: empirical formula CoI3